### v1

# Constrained inhomogeneous spherical equations: average-case hardness

Abstract.
In this paper we analyze computational properties of the Diophantine problem (and its search variant) for spherical equations $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ (and its variants) over the class of finite metabelian groups $G_{p,n}={\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$, where $n\in{\mathbb{N}}$ and $p$ is prime. We prove that the problem of finding solutions for certain constrained spherical equations is computationally hard on average (assuming that some lattice approximation problem is hard in the worst case).
Keywords. Spherical equations, finite groups, semidirect products, metabelian groups, average case complexity, hash function family, group-based cryptography.
2010 Mathematics Subject Classification. 20F16, 20F10, 68W30.
1. Introduction
The main goal of this paper is to create bridges between problems of computational group theory (namely the problem of finding a solution for a spherical equation) and assumptions of lattice-based cryptography.
Modern uses of lattices in the design of cryptographic primitives started in 1996 with the paper [3], where M. Ajtai introduced a class of random problems hard for probabilistic polynomial-time (PPT) algorithms on any cryptographically non-negligible sets of instances unless certain lattice approximation problems (such as ${\mathbf{SIVP}}_{\gamma}$, see Section 3.1) can be solved efficiently in the worst case (on every input). Since then Ajtai’s construction was studied extensively; it was improved in different ways, various interesting applications were found, and it gained a lot of popularity, see [37] for a survey. Nowadays, lattice-based cryptography appears to be the most promising branch of post-quantum cryptography.
1.1. Equations in groups
Let $F=F(Z)$ denote the free group on countably many generators $Z=\{z_{i}\}_{i=1}^{\infty}$. For a group $G$, an equation over $G$ with variables in $Z$ is an equality of the form $W=1$, where $W\in F\ast G$. If $W=z_{i_{1}}g_{1}\cdots z_{i_{k}}g_{k}$, with $z_{i_{j}}\in Z$ and $g_{j}\in G$, then we refer to $\{z_{i_{1}},\ldots,z_{i_{k}}\}$ as the set of variables and to $\{g_{1},\ldots,g_{k}\}$ as the set of constants (or coefficients) of $W$. We occasionally write $W(z_{1},\ldots,z_{k})$ or $W(z_{1},\ldots,z_{k};g_{1},\ldots,g_{k})$ to indicate that the variables in $W$ are precisely $z_{1},\ldots,z_{k}$ and (in the latter case) the constants are precisely $g_{1},\ldots,g_{k}$. A solution for an equation $W(z_{1},\ldots,z_{k})=1$ over $G$ is an assignment for variables $z_{1},\ldots,z_{k}$ that makes $W=1$ true.
In this paper we assume that $G$ comes equipped with a fixed generating set $X$ and elements of $G$ are given as products of elements of $X$ and their inverses. This naturally defines the length (or size) of the equation $W=1$ as the length of its left-hand side $W$.
The Diophantine problem over a group $G$ for a class of equations $C$ is an algorithmic question to decide whether a given equation $W=1$ in $C$ has a solution.
By definition, the Diophantine problem is a decision problem (a yes/no question). Additionally, one can study the corresponding search problem that requires to find a solution for $W=1$, provided one exists.
1.2. Constrained equations
A constrained equation over a group $G$ is an equation $W(z_{1},\dots,z_{n})=1$ equipped with a set $Z\subseteq G^{n}$. The Diophantine problem for a constrained equation requires to decide whether the following question has a solution or not:
$\left\{\begin{array}[]{l}W(z_{1},\dots,z_{n})=1,\\ (z_{1},\dots,z_{n})\in Z.\end{array}\right.$
In this paper we work with a set of the form $Z=Z_{1}\times\dots\times Z_{n}$, in which case every variable $z_{i}$ is constrained individually by $Z_{i}$.
Constrained equations in groups have received very little attention; only certain equations in free groups have been studied. For instance, in [9] V. Diekert demonstrated decidability and ${\mathbf{PSPACE}}$-hardness of the following problem: given an arbitrary equation
$w(x_{1},\dots,x_{n},a_{1},\dots,a_{m})=1$
in a free group $F_{m}$ and a list $H_{1},\dots,H_{n}$ of regular subsets of $F_{m}$, decide whether there exists a solution of this equation satisfying $x_{i}\in H_{i}$. V. Durnev showed in [11] that for a free group $F=F(a,b)$ there exists an equation of type
$w(x^{k},x_{1},\dots,x_{n})=[a,b]$
for which there is no algorithm to recognize if, for a given $k$, it has a solution satisfying $x_{1},\dots,x_{n}\in[F_{2},F_{2}]$.
1.3. Spherical equations
One class of equations over groups that has generated much interest is the class of quadratic equations: equations where each variable $z$ appears exactly twice (as either $z$ or $z^{-1}$). It was observed in the early 80’s [8, 38] that such equations have an affinity with the theory of compact surfaces (for instance, via their associated van Kampen diagrams). This geometric point of view sparked the initial interest in their study and has led to many interesting results, particularly in the realm of quadratic equations over free groups: solution sets were studied in [17], ${\mathbf{NP}}$-completeness was proved in [10, 21]. Systems of quadratic equations played an important role in the study of the first order theory of free groups (Tarski problem, [23]). quadratic equations in various classes of (infinite) groups such as hyperbolic groups (solution sets described in [18], ${\mathbf{NP}}$-complete by [22]), the first Grigorchuk group (decidability proved in [26], commutator width computed in [5]), free metabelian groups (${\mathbf{NP}}$-hard by [27], in ${\mathbf{NP}}$ for orientable equations by [28]), metabelian Baumslag–Solitar groups (${\mathbf{NP}}$-complete by [29]), etc.
We say that equations $W=1$ and $V=1$ are equivalent if there is an automorphism $\phi\in\operatorname{{Aut}}(F\ast G)$ such that $\phi$ is the identity on $G$ and $\phi(W)=V$. It is a well known consequence of the classification of compact surfaces that any quadratic equation over $G$ is equivalent, via an automorphism $\phi$, computable in time $O(|W|^{2})$, to an equation in exactly one of the following three standard forms (see [7, 17]):
(1) $\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $m\geq 1,$
(2) $\prod_{i=1}^{g}[x_{i},y_{i}]\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0,$
(3) $\prod_{i=1}^{g}x_{i}^{2}\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0.$
The number $g$ is the genus of the equation, and both $g$ and $m$ (the number of constants) are invariants. The standard forms are called, respectively, spherical, orientable of genus $g$, and non-orientable of genus $g$.
In this paper we investigate spherical equations in finite groups. We say that the equation (1) is a homogeneous form of a spherical equation. For $m\in{\mathbb{N}}$ define the set of all homogeneous equations with $m$ conjugates by
$\operatorname{Sph}_{m}=\left\{\,\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}=1\;\middle|\;c_{1},\dots,c_{m}\in G\,\right\}\mbox{ and }\operatorname{Sph}=\bigcup_{m=1}^{\infty}\operatorname{Sph}_{m}.$
It is easy to show that an equation of the form
(4) $\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}=c$
is equivalent to a spherical equation. We say that an equation (4) is an inhomogeneous form of a spherical equation.
Notice that spherical equations naturally generalize fundamental (Dehn) problems of group theory, as solving equations from $\operatorname{Sph}_{1}$ is the same as solving the word problem and solving equations from $\operatorname{Sph}_{2}$ is the same as solving the conjugacy problem. The complexity of solving equations in finite groups has been first studied by Goldmann and Russell [14] showing that the Diophantine problem in a fixed finite nilpotent group can be decided in polynomial time, while it is ${\mathbf{NP}}$-complete in every finite non-solvable group. For some recent results, see [12, 19, 30].
1.4. Groups under consideration
Consider the group ${\mathbb{Z}}_{p}^{n}=\left\{\,(x_{1},\ldots,x_{n})\;\middle|\;x_{i}\in{\mathbb{Z}}_{p}\,\right\}$. The group of units ${\mathbb{Z}}_{p}^{\ast}$ acts on ${\mathbb{Z}}_{p}^{n}$ by (scalar) multiplication
$(x_{1},\ldots,x_{n})\ \stackrel{{\scriptstyle\alpha}}{{\mapsto}}\ (\alpha x_{1},\ldots,\alpha x_{n}),$
where $\alpha\in{\mathbb{Z}}_{p}^{\ast}$ and $(x_{1},\ldots,x_{n})\in{\mathbb{Z}}_{p}^{n}$. The semidirect product $G={\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$ is a set of pairs $({\overline{x}},\alpha)$ equipped with the binary operation
$({\overline{x}},\alpha)({\overline{y}},\beta)=({\overline{x}}+\alpha{\overline{y}},\alpha\cdot\beta),$
with the identity $({\overline{0}},1)$. The following useful formulae are used throughout the paper without referencing:
$({\overline{x}},\alpha)^{-1}$ $=(-\alpha^{-1}{\overline{x}},\alpha^{-1}),$
$({\overline{x}},\alpha)^{-1}({\overline{y}},\beta)({\overline{x}},\alpha)$ $=(\alpha^{-1}((\beta-1){\overline{x}}+{\overline{y}}),\beta).$
1.5. Model of computation
We assume that all computations are performed on a random access machine. Elements of ${\mathbb{Z}}_{p}$ are given in binary as bit-strings of length $\lceil\log_{2}(p)\rceil$. Elements of ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$ are given as $n+1$-tuples of elements from ${\mathbb{Z}}_{p}$. If $f(n)$ is $O(T(n)\cdot n^{\varepsilon})$ for every $\varepsilon>0$, then we say that a function $f(n)$ is “nearly $T(n)$” and write $f(n)$ is $\tilde{O}(T(n))$. Operations in ${\mathbb{Z}}_{p}$ have the following complexity:
  • Addition and subtraction can be done in $O(\lceil\log_{2}(p)\rceil)$ time in a straightforward way.
  • Multiplication can be done in $\tilde{O}(\lceil\log_{2}(p)\rceil)$ time using fast Fourier transform.
  • Computing the multiplicative inverse of a unit modulo $p$ can be done in $O(\lceil\log_{2}(p)\rceil^{2})$ time using the extended Euclidean algorithm.
2. Preliminaries: knapsack problems in groups
Here we review several useful definitions of discrete optimization in groups. Let $G$ be a group generated by a finite set $X=\{x_{1},\ldots,x_{n}\}\subseteq G$. Elements in $G$ can be expressed as products of the generators in $X$ and their inverses. Hence, we can state the following combinatorial problem.
The subset sum problem ${\mathbf{SSP}}(G,X)$: Given $g_{1},\ldots,g_{k},g\in G$ decide if
(5) $g=g_{1}^{\varepsilon_{1}}\cdots g_{k}^{\varepsilon_{k}}$
for some $\varepsilon_{1},\ldots,\varepsilon_{k}\in\{0,1\}$.
By [31, Proposition 2.5], computational properties of ${\mathbf{SSP}}$ do not depend on the choice of a finite generating set $X$ and, hence, the problem can be abbreviated as ${\mathbf{SSP}}(G)$. The same paper, [31], provides a variety of examples of groups with ${\mathbf{NP}}$-complete (or polynomial time) subset sum problems.
Consider the infinitely generated group ${\mathbb{Z}}_{3}^{\omega}$ whose elements can be formally viewed as functions $f\colon{\mathbb{N}}\to{\mathbb{Z}}_{3}$ with finite support. For algorithmic purposes, we assume that elements of ${\mathbb{Z}}_{3}^{\omega}$ are encoded by finite ternary strings (as in [35, Section 4]). [36, Proposition 2.1] proves that ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{\omega})$ is ${\mathbf{NP}}$-complete, which can be reformulated as follows.
Proposition 2.1 ([36, Proposition 2.1]).
For every $m\geq 3$, ${\mathbf{SSP}}$ is ${\mathbf{NP}}$-complete for the class of finite groups $\{{\mathbb{Z}}_{m}^{n}\}_{n=1}^{\infty}$.
3. Preliminaries: lattice problems
3.1. Lattices
Here we review several definitions of the theory of lattices. Recall that a set of points $S\subseteq{\mathbb{R}}^{n}$ is discrete if every point $x\in{\mathbb{R}}^{n}$ has an $\varepsilon$-neighborhood that contains $x$ only. A lattice is a discrete subgroup of ${\mathbb{R}}^{n}$. An integer lattice is a subgroup of ${\mathbb{Z}}^{n}$. We discuss integer lattices only. We say that ${\mathcal{L}}\leq{\mathbb{Z}}^{n}$ is a full-rank lattice if the dimension of the corresponding vector space $\operatorname{{Span}}({\mathcal{L}})$ is $n$ (we can simply write $\dim({\mathcal{L}})=n$). We typically assume that ${\mathcal{L}}$ is a full-rank lattice.
The minimum distance of a lattice ${\mathcal{L}}\leq{\mathbb{R}}^{n}$ is the length of a shortest nonzero lattice vector
$\lambda_{1}({\mathcal{L}})=\min_{{\overline{v}}\in{\mathcal{L}}\setminus\{{\overline{0}}\}}\|{\overline{v}}\|.$
More generally, for $i=1,\ldots,n$, the $i$th successive minimum of ${\mathcal{L}}$ is
$\lambda_{i}({\mathcal{L}})=\min\left\{r\mid\dim(\operatorname{{Span}}(B(r)\cap{\mathcal{L}}))\geq i\right\},$
which is the minimum radius of a ball that contains at least $i$ linearly independent points. Below we recall several computational problems for lattices. Some of them are of direct importance to lattice-based cryptography (e.g., ${\mathbf{SIVP}}_{\gamma}$ and ${\mathbf{GapSVP}}_{\gamma}$) and others have more historical importance.
Shortest vector problem, ${\mathbf{SVP}}$. Given a basis of a lattice ${\mathcal{L}}\leq{\mathbb{Z}}^{n}$, find a shortest nonzero vector ${\overline{v}}\in{\mathcal{L}}$, i.e., a vector satisfying $\|{\overline{v}}\|=\lambda_{1}({\mathcal{L}})$.
Approximate shortest vector problem, ${\mathbf{SVP}}_{\gamma}$. Given a basis of a lattice ${\mathcal{L}}\leq{\mathbb{Z}}^{n}$, find a nonzero vector ${\overline{v}}\in{\mathcal{L}}$ satisfying $\|{\overline{v}}\|\leq\gamma\cdot\lambda_{1}({\mathcal{L}})$.
Decisional approximate SVP, ${\mathbf{GapSVP}}_{\gamma}$. Given a basis of a lattice ${\mathcal{L}}$, where either $\lambda_{1}({\mathcal{L}})\leq 1$ or $\lambda_{1}({\mathcal{L}})>\gamma$, decide which is the case.
Shortest independent vector problem, ${\mathbf{SIVP}}$. Given a basis of a full-rank lattice ${\mathcal{L}}\leq{\mathbb{Z}}^{n}$, find $n$ linearly independent vectors ${\overline{v}}_{1},\dots,{\overline{v}}_{n}\in{\mathcal{L}}$ satisfying $\max\|{\overline{v}}_{i}\|\leq\lambda_{n}({\mathcal{L}})$.
Approximate shortest independent vectors problem, ${\mathbf{SIVP}}_{\gamma}$. Given a basis of a full-rank lattice ${\mathcal{L}}\leq{\mathbb{Z}}^{n}$, find $n$ linearly independent vectors ${\overline{v}}_{1},\dots,{\overline{v}}_{n}\in{\mathcal{L}}$ satisfying
$\max\|{\overline{v}}_{i}\|\leq\gamma(n)\cdot\lambda_{n}({\mathcal{L}}).$
Formally, we say that ${\mathbf{SIVP}}_{\gamma}$ is hard in the worst case for probabilistic polynomial-time (PPT) algorithms, if for every PPT algorithm ${\mathcal{A}}$ and for every $n\in{\mathbb{N}}$ there is an $n$-dimensional basis $B_{n,{\mathcal{A}}}=\{{\overline{b}}_{1},\dots,{\overline{b}}_{n}\}\subseteq{\mathbb{Z}}^{n}$ satisfying
[…]
for every $d>0$, where the probability is taken over the coin-tosses of ${\mathcal{A}}$.
The above lattice problems have been intensively studied and appear to be intractable, except for very large approximation factors $\gamma(n)$. Known polynomial-time algorithms like the Lenstra–Lenstra–Lovász [25] and its descendants obtain only slightly subexponential approximation factors for all the above problems. Known algorithms that obtain polynomial approximation factors, such as [20, 4, 34, 1], either require superexponential time, or exponential time and space.
Many lattice problems are ${\mathbf{NP}}$-hard, even to approximate to within various sub-polynomial $n^{o(1)}$ approximation factors. However, such hardness is not of any direct consequence to cryptography, since lattice-based cryptographic constructions so far rely on polynomial approximation problems factors $\gamma(n)\geq n$. Indeed, there is evidence that for factors $\gamma(n)\geq\sqrt{n}$, the lattice problems relevant to cryptography are not ${\mathbf{NP}}$-hard, because they lie in ${\mathbf{NP}}\cap{\mathbf{coNP}}$ [15, 2]. For a survey on lattice-based cryptography see [37].
3.2. Short integer solution problem
In general, the short integer solution problem can be formulated as follows.
Short integer solution (${\mathbf{SIS}}$) problem. For a given matrix $A\in{\mathbb{Z}}_{p}^{n\times m}$, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ satisfying
(7) $\left\{\begin{array}[]{l}A{\overline{x}}\equiv_{p}{\overline{0}},\\ {\overline{x}}\mbox{ is short and nontrivial},\\ \end{array}\right.$
assuming that a solution exists.
In other words, ${\mathbf{SIS}}$ requires to find a short and nontrivial solution for a homogeneous system of linear congruences $A{\overline{x}}\equiv_{p}{\overline{0}}$. In the original paper [3], ${\overline{x}}$ was called short if $\|{\overline{x}}\|_{2}\leq n$ (can be relaxed to $\|{\overline{x}}\|_{2}\leq poly(n)$). In [16], instead of $\|{\overline{x}}\|_{2}\leq n$, the following two constraints on ${\overline{x}}$ were discussed:
  • ${\overline{x}}\neq 0$ and $x_{i}\in\{0,1\}$;
  • ${\overline{x}}\neq 0$ and $x_{i}\in\{-1,0,1\}$.
We denote the corresponding versions of ${\mathbf{SIS}}$ by ${\mathbf{SIS}}_{\{0,1\}}$ and ${\mathbf{SIS}}_{\{-1,0,1\}}$. To discuss the average case complexity of ${\mathbf{SIS}}$ we use the uniform distribution on its instances.
Randomized ${\mathbf{SIS}}$ problem. For a matrix $A\leftarrow{\mathbb{Z}}_{p}^{n\times m}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{0}}$, and ${\overline{x}}$ is short and nontrivial.
Theorem 3.1 (Goldreich, Goldwasser, Halevi, [16, Theorem 1], cf. Ajtai, [3, Theorem 1]).
Suppose that a PPT algorithm ${\mathcal{A}}$ solves the randomized ${\mathbf{SIS}}_{\{0,1\}}$ problem (or ${\mathbf{SIS}}_{\{-1,0,1\}}$ problem) with parameters $n,m,p$ satisfying
(8) $n\log(p)<m<\tfrac{p}{2n^{4}}\ \mbox{ and }\ p=O(n^{c})\mbox{ for some }c>0,$
with probability at least $n^{-c_{0}}$ for some fixed constant $c_{0}>0$, where the probability is taken over the choice of the instance as well as the coin-tosses of ${\mathcal{A}}$. Then there is a PPT algorithm that solves ${\mathbf{GapSVP}}_{\gamma=pn^{6}}$ and ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ (among others) on every $n$-dimensional lattice with probability at least $1-2^{-n}$.  [^1] For better bounds and approximation factors of $\gamma$ see [33].
In a similar way we can define deterministic and randomized inhomogeneous short integer solution (ISIS) problem.
Randomized ${\mathbf{ISIS}}$ problem. For a matrix $A\leftarrow{\mathbb{Z}}_{p}^{n\times m}$ and a vector ${\overline{y}}\in{\mathbb{Z}}_{p}^{n}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{y}}$, and ${\overline{x}}$ is short.
Corollary 3.2.
The statement of Theorem 3.1 holds for the randomized ${\mathbf{ISIS}}_{\{0,1\}}$ problem.
Proof.
Indeed, suppose that there is a PPT algorithm ${\mathcal{A}}$ that solves ${\mathbf{ISIS}}_{\{0,1\}}$ problem with probability at least $n^{-c_{0}}$. Then for an instance $A$ of ${\mathbf{SIS}}_{\{0,1\}}$ with columns ${\overline{v}}_{1},\dots,{\overline{v}}_{m}$
  1. (1)
    “guess” a nonzero index $i$ in a solution ${\overline{x}}\in\{0,1\}^{m}$,
  2. (2)
    form an instance $(A^{\prime},-{\overline{v}}_{i})$, where $A^{\prime}$ is obtained by deleting the $i$th column from $A$,
  3. (3)
    solve $(A^{\prime},-{\overline{v}}_{i})$ using ${\mathcal{A}}$.
This gives a PPT algorithm that solves ${\mathbf{SIS}}_{\{0,1\}}$ with probability at least $n^{-c_{0}}$. ∎
4. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
In this section we analyse complexity of solving a spherical equation (1) over $G_{p,n}$, with constants $c_{i}=({\overline{c}}_{i},\beta_{i})\in G_{p,n}$ and unknowns $z_{1},\dots,z_{m}$.
Lemma 4.1.
[…]
  • (S1)
    $\beta_{1}\cdots\beta_{m}=1$;
  • (S2)
    $\sum_{i=1}^{m}B_{i}\alpha_{i}^{-1}((\beta_{i}-1){\overline{z}}_{i}+{\overline{c}}_{i})\equiv_{p}{\overline{0}}$, where $B_{i}=\beta_{1}\cdots\beta_{i-1}$.
Proof.
Straightforward verification. ∎
4.1. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$: generic-case
Here, we investigate generic-case hardness for spherical equations, i.e., hardness of a “typical” equation, see [32, Chapter 10] for basic definitions of generic-case complexity.
Proposition 4.2.
If $\beta_{i}\neq 1$ for some $i$, then (1) has a solution and a solution can be found in polynomial time.
Proof.
It takes nearly linear time to compute $B_{1},\dots,B_{m}$ and check the condition (S1). If $\beta_{i}\neq 1$, then the following assignment:
  • $\alpha_{1}=\dots=\alpha_{m}=1$,
  • ${\overline{z}}_{j}={\overline{0}}$ for $j\neq i$,
  • ${\overline{z}}_{i}\ \ \equiv_{p}\ \ \tfrac{1}{B_{i}(\beta_{i}-1)}\sum_{j=1}^{m}B_{j}{\overline{c}}_{j}$
is a solution that can be computed in polynomial time. ∎
Next, we claim that the property $\exists i\ \beta_{i}\neq 1$ is strongly generic, i.e., a typical equation satisfies this property. Since the problem involves three parameters, $n,m$, and $p$, we use the following stratification for the set of instances of the uniform problem. For $s\in{\mathbb{N}}$ define a set of pairs
${\mathcal{I}}_{s}=\left\{\,(E_{m},G_{p,n})\;\middle|\;E_{m}\in\operatorname{Sph}_{m}(G_{p,n}),\ m,n,p\leq s\,\right\}$
equipped with the uniform distribution. Then the following holds:
$|G_{p,n}|$ $=p^{n}\cdot(p-1),$
$|\operatorname{Sph}_{m}|$ $=(p^{n}\cdot(p-1))^{m},$
$|{\mathcal{I}}_{s}|$ $=\sum_{p,n,m\leq s}(p^{n}\cdot(p-1))^{m}.$
Lemma 4.3.
$\Pr\left\{\,(E_{m},G_{p,n})\in{\mathcal{I}}_{s}\;\middle|\;m\geq s/2\,\right\}\to 1$ exponentially fast as $s\to\infty$.
Proof.
For any fixed $p\geq 2,n\geq 1$ we have
$V_{s}=\sum_{m=0}^{s}(p^{n}\cdot(p-1))^{m}=\frac{(p^{n}\cdot(p-1))^{s+1}-1}{p^{n}\cdot(p-1)-1},$
which implies that
$\frac{V_{s/2}}{V_{s}}=\frac{(p^{n}\cdot(p-1))^{s/2+1}-1}{(p^{n}\cdot(p-1))^{s+1}-1}\leq\frac{1}{(p^{n}\cdot(p-1))^{s/2}}\leq\frac{1}{2^{s/2}}$
that converges to 00 exponentially fast. The obtained bound $\frac{1}{2^{s/2}}$ does not depend on $p$ or $n$. Therefore, the bound works on the whole ${\mathcal{I}}_{s}$. ∎
Lemma 4.4.
$\Pr\left\{\,(E_{m},G_{p,n})\in{\mathcal{I}}_{s}\;\middle|\;\exists i\ \beta_{i}\neq 1\,\right\}\to 1$ exponentially fast as $s\to\infty$.
Proof.
By Lemma 4.3, we may assume that $m\geq s/2$, in which case
$\Pr\left\{\,(E_{m},G_{p,n})\in{\mathcal{I}}_{s}\;\middle|\;\exists i\ \beta_{i}\neq 1\,\right\}\geq 1-\frac{1}{(p-1)^{s/2}},$
which converges to $1$ exponentially fast as $s\to\infty$. ∎
Corollary 4.5.
The Diophantine (decision) problem for spherical equations over $\{G_{p,n}\}$ is decidable in strongly generically linear time. The corresponding search problem can be solved in strongly generically polynomial time.
Proof.
Follows from Proposition 4.2 and Lemma 4.4
Since spherical equations in which at least one coefficient $c_{i}=({\overline{c}}_{i},\beta_{i})$ satisfies $\beta_{i}\neq 1$ are computationally easy, we put a restriction on the coefficients that we use in spherical equations. Define the set
(9) $C_{p,n}=\left\{\,({\overline{c}},1)\;\middle|\;{\overline{c}}\in{\mathbb{Z}}_{p}^{n}\,\right\}.$
From now on we assume that all $c_{i}\in C_{p,n}$. In that case condition (S1) of Lemma 4.1 is trivially satisfied and condition (S2) translates into the following:
(10) $\exists\alpha_{i}\in{\mathbb{Z}}_{p}^{\ast}\ \ \mbox{ s.t. }\ \sum_{i=1}^{m}\alpha_{i}^{-1}{\overline{y}}_{i}\equiv_{p}{\overline{0}}.$
The obtained condition defines a homogeneous system of linear congruences that does not allow zero values for unknowns $\alpha_{i}$, which makes the problem nontrivial (we conjecture it is ${\mathbf{NP}}$-hard).
4.2. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$: worst case
Here we investigate the worst case complexity for spherical equations. For ${\overline{w}}\in{\mathbb{Z}}_{3}^{n}$ define
$c_{\overline{w}}=({\overline{w}},1)\in{\mathbb{Z}}_{3}^{n}\rtimes{\mathbb{Z}}_{3}^{\ast}.$
Consider an instance $I$ of ${\overline{v}}_{1},\dots,{\overline{v}}_{m},{\overline{v}}\in{\mathbb{Z}}_{3}^{n}$ of ${\mathbf{SSP}}$. Recall that $I$ is a positive instance if $\varepsilon_{1}{\overline{v}}_{1}+\dots+\varepsilon_{m}{\overline{v}}_{m}={\overline{v}}$ for some $\varepsilon_{1},\dots,\varepsilon_{m}\in\{0,1\}$. For $I$ define ${\overline{v}}^{\prime}={\overline{v}}+\sum{\overline{v}}_{i}$ and the spherical equation $E_{I}$ as
(11) $\prod_{i}^{m}z_{i}^{-1}c_{{\overline{v}}_{i}}z_{i}=c_{{\overline{v}}^{\prime}}.$
Proposition 4.6.
$I$ is a positive instance of ${\mathbf{SSP}}$Leftrightarrow\ \ $E_{I}$ has a solution.
Proof.
By design we have
$I\mbox{ is a positive instance of }{\mathbf{SSP}}$ $\ \Leftrightarrow\ {\overline{v}}=\sum_{i=1}^{m}\varepsilon_{i}{\overline{v}}_{i}\mbox{ for }\varepsilon_{1},\dots,\varepsilon_{m}\in\{0,1\}$
$\ \Leftrightarrow\ {\overline{v}}+\sum_{i=1}^{m}{\overline{v}}_{i}=\sum_{i=1}^{m}\beta_{i}{\overline{v}}_{i}\mbox{ for }\beta_{1},\dots,\beta_{m}\in\{1,2\}={\mathbb{Z}}_{3}^{\ast}$
$\ \Leftrightarrow\ {\overline{v}}+\sum_{i=1}^{m}{\overline{v}}_{i}=\sum_{i=1}^{m}\alpha_{i}^{-1}{\overline{v}}_{i}\mbox{ for }\alpha_{1},\dots,\alpha_{m}\in\{1,2\}={\mathbb{Z}}_{3}^{\ast}$
[…]
Corollary 4.7.
The Diophantine problem for spherical equations over ${\mathbb{Z}}_{3,n}\rtimes{\mathbb{Z}}_{3}^{\ast}$ is ${\mathbf{NP}}$-complete.
Proof.
By Proposition 4.6, $I\to E_{I}$ is a many-one polynomial-time reduction of ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{n})$ to spherical equations over ${\mathbf{SSP}}({\mathbb{Z}}_{3,n}\rtimes{\mathbb{Z}}_{3}^{\ast})$. By Proposition 2.1, ${\mathbf{SSP}}$ is ${\mathbf{NP}}$-complete for groups $\{{\mathbb{Z}}_{3}^{n}\}_{n=1}^{\infty}$. Hence the result. ∎
5. Constrained spherical equations: average-case hardness
In this section we discuss spherical equations over the groups $G_{p,n}$ with constraints on values of $z_{i}$’s. In full generality the problem can be formulated as follows. Given a group $G\in{\mathcal{G}}$, a spherical equation $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ over $G$, and subsets $Z_{1},\dots,Z_{m}\subseteq G$, find $z_{1},\dots,z_{m}\in G$ satisfying
$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1\\ z_{i}\in Z_{i}\end{array}\right.$
or a similar inhomogeneous form
$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\\ z_{i}\in Z_{i}.\end{array}\right.$
The problem to decide if a given constrained inhomogeneous spherical equation has a solution is abbreviated ${\mathbf{CISE}}$.
It is not hard to prove that ${\mathbf{CISE}}$ for groups $G_{p,n}$ is hard in the worst case. In fact, for $p=3$, the proof of Proposition 4.6 works as is, with constraints $Z_{i}=\{({\overline{0}},1),({\overline{0}},2)\}$ (so chosen sets $Z_{i}$ effectively define no constraints). Using these constraints we can easily achieve the same result for any odd value of $p$.
Theorem 5.1.
${\mathbf{CISE}}$ is ${\mathbf{NP}}$-hard for groups $\{G_{p,n}\}_{n}$ for any fixed odd prime $p$.
Proof.
For a given instance ${\overline{v}}_{1},\dots,{\overline{v}}_{m},{\overline{v}}\in{\mathbb{Z}}_{p}^{n}$, call it $I$, of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$, as in the proof of Proposition 4.6, construct a constrained inhomogeneous spherical equation $E_{I}$ over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{{\overline{v}}_{i}}z_{i}=c_{{\overline{v}}^{\prime}}\\ z_{i}=({\overline{0}},1)\mbox{ or }({\overline{0}},2^{-1})\end{array}\right.$
and observe that $I$ is a positive instance of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$ if and only if $E_{I}$ has a solution. This gives a polynomial time reduction from ${\mathbf{SSP}}$ over groups $\{{\mathbb{Z}}_{p}^{n}\}_{n=1}^{\infty}$ to decidability of constrained inhomogeneous spherical equations over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$. ∎
Our next goal is to demonstrate the average-case hardness of constrained spherical equations. To achieve this, we randomize equations as follows. First, we stratify parameters:
  • the parameter $n\in{\mathbb{N}}$ is the main independent parameter;
  • parameters $m$ and $p$ are functions of $n$ satisfying conditions (8).
Then, for an arbitrary but fixed $n\in{\mathbb{N}}$, consider the set of coefficients $C_{p,n}$ defined in (9) and a (finite) set of constrained inhomogeneous spherical equations
(12) $\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{n}\\ z_{i}=({\overline{0}},1)\mbox{ or }({\overline{0}},2^{-1}).\end{array}\right.$
The problem to find a solution of (12) is called ${\mathbf{CISE}}_{\{1,2\}}$. Also, we want to study slightly relaxed equations
(13) $\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{n}\\ z_{i}=({\overline{0}},1),({\overline{0}},2^{-1}),\mbox{ or }({\overline{0}},3^{-1}).\end{array}\right.$
The problem to find a solution of (13) is called ${\mathbf{CISE}}_{\{1,2,3\}}$.
Randomized constrained inhomogeneous spherical equation problem. Find a solution for a uniformly distributed system (12) (or (13)).
Theorem 5.2.
Suppose that a PPT algorithm ${\mathcal{A}}$ solves the randomized ${\mathbf{CISE}}_{\{1,2\}}$ problem (or ${\mathbf{CISE}}_{\{1,2,3\}}$ problem) with parameters $n,m,p$ satisfying (8) with probability at least $n^{-c_{0}}$ for some fixed constant $c_{0}>0$, where the probability is taken over the choice of the instance as well as the coin-tosses of ${\mathcal{A}}$. Then there is a PPT algorithm that solves ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ for any lattice with overwhelming probability (e.g. $1-2^{-n}$).
Proof.
Consider an arbitrary instance of $(A,{\overline{y}})$ of ${\mathbf{ISIS}}_{\{0,1\}}$. Let ${\overline{v}}_{1},\dots,{\overline{v}}_{m}$ be the system of columns of $A$. Consider the equation (11), where, as in Section 4.2, $c_{\overline{v}}=({\overline{v}},1)$ and ${\overline{v}}^{\prime}={\overline{y}}+\sum_{i=1}^{m}{\overline{v}}_{i}$. Then for ${\overline{x}}=(x_{1},\dots,x_{m})\in\{0,1\}^{m}$ we have
${\overline{x}}\mbox{ is a solution for $(A,{\overline{y}})$}$ $\ \ \Leftrightarrow\ \ \sum_{i=1}^{m}x_{i}{\overline{v}}_{i}={\overline{y}}$
$\ \ \Leftrightarrow\ \ \sum_{i=1}^{m}(x_{i}+1){\overline{v}}_{i}={\overline{y}}+\sum_{i=1}^{m}{\overline{v}}_{i}$
$\ \ \Leftrightarrow\ \ \prod_{i=1}^{m}({\overline{0}},(x_{i}+1)^{-1})^{-1}c_{{\overline{v}}_{i}}({\overline{0}},(x_{i}+1)^{-1})=c_{{\overline{v}}^{\prime}}$
[…]
This establishes a one-to-one correspondence between ${\mathbf{ISIS}}_{\{0,1\}}$ and ${\mathbf{CISE}}_{\{1,2\}}$ that translates the uniform distribution for the instances of ${\mathbf{ISIS}}_{\{0,1\}}$ to the uniform distribution for the instances of ${\mathbf{CISE}}_{\{1,2\}}$ (the element ${\overline{v}}^{\prime}\in C_{n}$ is uniformly distributed because ${\overline{y}}$ uniformly distributed). A similar reduction can be used to establish a one-to-one correspondence between ${\mathbf{SIS}}_{\{-1,0,1\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$. ∎
Remark 1.
The one-to-one correspondence established in the proof of Theorem 5.2 also implies that uniformly chosen instances of ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$ have solutions with probability approaching $1$ as $n\to\infty$.
Remark 2.
Similar results can be proven for the class of symmetric groups $\{S_{n}\}_{n=1}^{\infty}$, or any other class of groups that contain the groups $G_{p,n}$ in a “compact” way.
6. Spherical functions
Let $G$ be a group. For ${\overline{c}}=(c_{1},\dots,c_{m})\in G^{m}$ define a function $f_{\overline{c}}\colon G^{m}\to G$ by
$(z_{1},\dots,z_{m})\ \ \stackrel{{\scriptstyle f_{\overline{c}}}}{{\mapsto}}\ \ (z_{1}^{-1}c_{1}z_{1})\cdots(z_{m}^{-1}c_{m}z_{m}).$
We call $f_{\overline{c}}$ a spherical function because the problem of finding an $f_{\overline{c}}$-preimage of $g\in G$ is equivalent to finding a solution for a spherical equation $(z_{1}^{-1}c_{1}z_{1})\cdots(z_{m}^{-1}c_{m}z_{m})=g$. Fix distinct $g_{0},g_{1}\in G$ and define a function $H_{\overline{c}}\colon\{0,1\}^{m}\to G$ by
$(b_{1},\dots,b_{m})\ \ \stackrel{{\scriptstyle H_{\overline{c}}}}{{\mapsto}}\ \ (g_{b_{1}}^{-1}c_{1}g_{b_{1}})\cdots(g_{b_{m}}^{-1}c_{m}g_{b_{m}}).$
called a $0/1$-spherical function of length $m$. Similarly, we can define $-1/0/1$-spherical functions of length $m$.
Consider the family of groups $\{G_{p,n}\}$ and $g_{0}=({\overline{0}},1),g_{1}=({\overline{0}},2^{-1})\in G_{p,n}$, where $2^{-1}=\tfrac{1}{2}(p+1)$ is the multiplicative inverse of $2$ modulo an odd prime $p$. As in Section 5, $n$ is considered to be the main parameter; the parameters $m$ and $p$ are functions of $n$ satisfying conditions (8). Define a system of functions
${\mathcal{H}}_{n}=\left\{\,H_{\overline{c}}\;\middle|\;c_{1},\dots,c_{m}\in C_{p,n}\,\right\}\mbox{ and }{\mathcal{H}}=\bigcup_{n=1}^{\infty}{\mathcal{H}}_{n}.$
Proposition 6.1.
If ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ is hard in the worst case, then ${\mathcal{H}}$ is a one-way function family.
Proof.
To prove that ${\mathcal{H}}$ is a one-way function family it is sufficient to show that for every PPT algorithm ${\mathcal{A}}$ the sequence of values
$P_{\mathcal{A}}(n)\ =\ \Pr[{\mathcal{A}}(1^{n},{\overline{c}},H_{\overline{c}}({\overline{x}}))\in H_{\overline{c}}^{-1}(H_{\overline{c}}({\overline{x}}))]$
converges to 00 faster than every $1/poly(n)$, where the probability is taken over
  • uniform choices of ${\overline{c}}\in(C_{p,n})^{m}$,
  • uniform choices of ${\overline{x}}\in\{0,1\}^{n}$, and
  • the coin-tosses of ${\mathcal{A}}$.
It is not difficult to check that for uniformly distributed ${\overline{c}}$ and ${\overline{x}}$, the values of $H_{\overline{c}}({\overline{x}})\in C_{p,n}$ are distributed nearly uniformly (the sequence of distributions on $C_{p,n}$ converges to the uniform distribution exponentially fast in terms of $n$) and we may assume that ${\mathcal{A}}$ deals with uniformly randomized ${\mathbf{CISE}}_{\{1,2\}}$.
Therefore, if it is not true that $P_{\mathcal{A}}(n)$ is $o(n^{-d})$ for some PPT algorithm ${\mathcal{A}}$ and some $d>0$, then $P_{\mathcal{A}}(n)\geq c\cdot n^{-d}$ (for some $c>0$) satisfied for infinitely many indices $n$. By Theorem 5.2 that means that there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6). ∎
Proposition 6.2.
If ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ is hard in the worst case, then ${\mathcal{H}}$ is a collision-free function family.
Proof.
To prove that ${\mathcal{H}}$ is a collision-free hash function family it is sufficient to show that for every PPT algorithm ${\mathcal{A}}$ the sequence of values
$P_{\mathcal{A}}(n)\ =\ \Pr[({\overline{x}},{\overline{y}})={\mathcal{A}}(1^{n},{\overline{c}})\ \&\ H_{\overline{c}}({\overline{x}})=H_{\overline{c}}({\overline{y}}))]$
converges to 00 faster than every $1/poly(n)$, where the probability is taken over
  • uniform choices of ${\overline{c}}\in(C_{p,n})^{m}$,
  • the coin tosses of ${\mathcal{A}}$.
If this condition is not satisfied, then there is a PPT algorithm ${\mathcal{A}}$ and $c,d>0$ satisfying $P_{\mathcal{A}}(n)\geq c\cdot n^{-d}$ for infinitely many indices $n$. Observe that
$H_{\overline{c}}({\overline{x}})=H_{\overline{c}}({\overline{y}})$ $\ \ \Leftrightarrow\ \ \sum_{j=1}^{m}x_{j}{\overline{c}}_{j}=\sum_{j=1}^{m}y_{j}{\overline{c}}_{j}$
$\ \ \Leftrightarrow\ \ \sum_{j=1}^{m}(2+x_{j}-y_{j}){\overline{c}}_{j}={\overline{0}},\ \ \mbox{where }2+x_{j}-y_{j}\in\{1,2,3\}$
[…]
Thus, if we can efficiently find collisions for uniformly distributed $0/1$-spherical equations $H_{\overline{c}}\in{\mathcal{H}}_{n}$ for infinitely many indices $n$, then we can efficiently find solutions for uniformly distributed ${\mathbf{CISE}}_{\{1,2,3\}}$. Then, by Theorem 5.2, there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6). Contradiction. ∎
7. The acyclic graph word problem: average-case hardness
${\mathbf{CISE}}$ can be naturally reduced to different knapsack-type problems in groups. One such reduction is discussed in this section. The acyclic graph word problem was introduced in [13] as a convenient generalization of ${\mathbf{SSP}}$ and as a tool for studying ${\mathbf{SSP}}$. Let $X$ be a generating set for $G$.
The acyclic graph word problem, ${\mathbf{AGWP}}(G,X)$: Given an acyclic directed graph $\Gamma$ with edges labeled by letters in $X\cup X^{-1}\cup\{\varepsilon\}$ with two marked vertices, $\alpha$ and $\omega$, decide whether there is an oriented path in $\Gamma$ from $\alpha$ to $\omega$ labeled by a word $w$ such that $w=1$ in $G$.
It is easy to show that complexity of ${\mathbf{AGWP}}$ does not depend on a choice of a generating set for $G$ and the problem can be abbreviated ${\mathbf{AGWP}}(G)$. Also, we can work with edges labeled with words over the alphabet $X$. There is a natural polynomial-time reduction from (decision/search)-${\mathbf{CISE}}$ to (decision/search)-${\mathbf{AGWP}}$. Indeed, for an instance of ${\mathbf{CISE}}$
$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\\ z_{i}\in Z_{i},\end{array}\right.$
we can construct a labelled digraph $\Gamma=(V,E)$, where
  • $V=\{0,1,\dots,m,m+1\}$
  • $E=\left\{\,j-1\stackrel{{\scriptstyle z^{-1}c_{j}z}}{{\longrightarrow}}j\;\middle|\;z\in Z_{j},\ j=1,\dots,m\,\right\}\cup\{m\stackrel{{\scriptstyle c^{-1}}}{{\longrightarrow}}m+1\}$
  • the starting point $\alpha$ is 00,
  • the terminal point $\omega$ is $m+1$.
The construction is visualized in Figure 1. By construction of $\Gamma$, the following statement holds.
Figure 1. An instance of ${\mathbf{AGWP}}$ corresponding to an instance of ${\mathbf{CISE}}$ (assuming that $Z_{j}=\{g_{1},g_{2}\}$).
Proposition 7.1.
The instance of ${\mathbf{CISE}}$ is positive if and only if the instance of ${\mathbf{AGWP}}$ is positive.
The reduction described above induces a system of probability measures $\{\mu_{n}\}_{n\in{\mathbb{N}}}$ on instances of ${\mathbf{AGWP}}$ over the class of groups ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$. In fact, each $\mu_{n}$ gives uniform distribution on a subset of instances of ${\mathbf{AGWP}}$ of type shown in Figure 7.1. That establishes a one-to one correspondence between uniformly randomized ${\mathbf{CISE}}_{\{1,2\}}$ described in Section 5 and ${\mathbf{AGWP}}$ endowed with $\{\mu_{n}\}_{n\in{\mathbb{N}}}$, which implies average-case hardness of ${\mathbf{AGWP}}$ (again, assuming (6)).
The converse reduction appears to be implausible. In fact, for some groups the converse is clearly not true, e.g., for ${\mathbb{Z}}_{3}^{\omega}$ (and for the class of groups $\{{\mathbb{Z}}_{3}^{n}\}_{n\in{\mathbb{N}}}$) we have
  • ${\mathbf{CISE}}$ is polynomial-time decidable, but
  • ${\mathbf{SSP}}$ (and hence ${\mathbf{AGWP}}$) is ${\mathbf{NP}}$-complete.
8. Open problems and questions for further study
In this section we outline several problems related to spherical equations, constrained spherical equations, relations between constrained/unconstrained problems, average-case complexity, and design of hash functions.
8.1. Spherical equations over finite groups: constrained versus unconstrained
Problem 8.1.
Does there exist a class of finite groups ${\mathcal{G}}$ satisfying the following conditions:
  • the Diophantine problem for spherical equations over ${\mathcal{G}}$ is efficiently decidable,
  • the Diophantine problem for constrained spherical equations is computationally hard?
Denote by $\mathbb{P}$ the set of all prime numbers. It was proved in [30] that the Diophantine problem for spherical equations over ${\mathcal{G}}=\{\operatorname{{GL}}(2,p)\}_{p\in\mathbb{P}}$ can be solved efficiently.
Problem 8.2.
Does $\{\operatorname{{GL}}(2,p)\}_{p\in\mathbb{P}}$ satisfy the second condition of Problem 8.1?
Problem 8.2 can be generalized as follows.
Problem 8.3.
Does any family of classical finite groups (such as $\operatorname{{SL}}(n,p)$, $\operatorname{{PSL}}(n,p)$, etc., see [24]) satisfy conditions of Problem 8.1?
8.2. Constrained spherical equations: foundation of average-case hardness
In Section 5 we showed that for groups $\{G_{p,n}\}$ constrained spherical equations (randomized in a certain way) are hard on average by establishing a one-to-one correspondence between ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{ISIS}}_{\{0,1\}}$. Average-case hardness of ${\mathbf{ISIS}}_{\{0,1\}}$ is linked (by M. Ajtai) to the worst-case hardness of lattice approximation problems, such as ${\mathbf{SIVP}}_{\gamma}$, which makes hardness of ${\mathbf{SIVP}}_{\gamma}$ the foundation of average hardness for ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{ISIS}}_{\{0,1\}}$. Informally, is it possible to untie ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{ISIS}}_{\{0,1\}}$ from ${\mathbf{SIVP}}_{\gamma}$?
One way to approach this question is to utilize a discrete logarithm type self-reduction idea. For ${\mathbf{ISIS}}_{\{0,1\}}$ that means to design a class of randomized self-reductions $\Phi$ between instances of ${\mathbf{ISIS}}_{\{0,1\}}$ and use $\Phi$ to enhance ${\mathbf{ISIS}}$-solvers as follows. If $A$ is an ${\mathbf{ISIS}}$-solver, then an enhanced ${\mathbf{ISIS}}$-solver $B$ for a given instance $I$ performs the following:
  • (a)
    Apply $A$ to $I$ and if it succeeds, then output the result.
  • (b)
    Choose a random $\varphi\in\Phi$ and apply $A$ to $\varphi(I)$. If it succeeds and produces a solution for $\varphi(I)$, then use it to “reconstruct” and output a solution for $I$.
  • (c)
    Repeat (b) until it succeeds.
If $\Phi$ is sufficiently large, then this approach might work. Notice that existence of uniform self-reductions for ${\mathbf{ISIS}}$ (like for the discrete logarithm problem) appears to be implausible.
Problem 8.4.
Design a class of randomized self-reductions $\Phi$ for ${\mathbf{ISIS}}_{\{0,1\}}$ and use them to prove the following. If ${\mathbf{ISIS}}_{\{0,1\}}$ can be solved by a PPT algorithm $A$ on a non-negligible set of instances, then it can be solved by a PPT algorithm $B$ on every instance with overwhelming probability.
Similarly, it would be interesting to investigate self-reductions for ${\mathbf{CISE}}$ over some groups ${\mathcal{G}}$ and use them to demonstrate that the worst-case hardness for ${\mathbf{CISE}}$ implies its average-case hardness.
Problem 8.5.
For a class of spherical equations over a class of groups ${\mathcal{G}}$, design a class of randomized self-reductions for ${\mathbf{CISE}}$ and use them to prove the following. If ${\mathbf{CISE}}$ can be solved by a PPT algorithm $A$ on a non-negligible set of instances, then it can be solved by a PPT algorithm $B$ on every instance with overwhelming probability.
8.3. Constrained spherical equations: generalization
As we have shown in Section 7, instances of ${\mathbf{CISE}}$ can be transformed into instances of ${\mathbf{AGWP}}$ of certain type (shown in Figure 7.1). Let us generalize the digraph shown in Figure 7.1 and modify it into a finite-state transducer. For a table of elements $C=\{c_{ij}\}$ from $G$ (where $i=1,\dots,m$ and $j=0,1$) define a transducer $\Gamma_{C}$ shown in Figure 2.
Figure 2. The transducer $\Gamma_{C}$.
Each $\Gamma_{C}$ defines a function $J_{C}\colon\{0,1\}^{m}\to G$ as follows:
$(b_{1},\dots,b_{m})\ \ \stackrel{{\scriptstyle J_{C}}}{{\mapsto}}\ \ c_{1b_{1}}\cdots c_{mb_{m}}.$
By design, the problem of finding a $J_{C}$-preimage $(b_{1},\dots,b_{m})\in\{0,1\}^{m}$ of a given element $c$ generalizes ${\mathbf{CISE}}$. Also, notice that the functions $J_{C}$ generalize hash function construction introduced in [40] and its numerous variations (e.g., [39, 6]), including those utilizing monoids.
Problem 8.6.
Investigate average-case hardness of computing $J_{C}^{-1}$.
The most promising approach to Problem 8.6 appears to be via self-reductions similar to those discussed in Section 8.2. Furthermore, it might be easier to design self-reductions for functions $J_{C}$ because $J_{C}$ does not require pairs $c_{i0}$ and $c_{i1}$ to be conjugate.
Problem 8.7.
For a class of groups ${\mathcal{G}}=\{G_{n}\}$ design a randomized self-reduction for the problem of computing $J_{C}^{-1}$ satisfying the following: if $J_{C}^{-1}$ can be computed by a PPT algorithm on a non-negligible set of instances, then it can be computed by a PPT algorithm on every instance with overwhelming probability.
Problem 8.8.
For a class of groups ${\mathcal{G}}$ investigate cryptographic properties for the family of functions $\{J_{C}\}_{C\in G^{2\times m}}$.
8.4. Constrained spherical equations: learning without errors
Consider a black-box device $D$ that computes values of a “hidden” function $H_{\overline{c}}\colon\{0,1\}^{m}\to G$ for some ${\overline{c}}\in G^{m}$ hardwired inside of $D$. The device has a single button which, when pressed, produces a pair $({\overline{b}},g)$, where
  • ${\overline{b}}=(b_{1},\dots,b_{m})\in\{0,1\}^{m}$ is chosen uniformly randomly,
  • $g=H_{\overline{c}}({\overline{b}})$.
To learn the hidden function means to find the secret element ${\overline{c}}$ using a number of sampled pairs $({\overline{b}},g)$.
Problem 8.9.
Investigate computational complexity of learning ${\overline{c}}$ from a hidden function $H_{\overline{c}}$ for different (classes of) finite/infinite groups.
In a similar way, we can consider a device $D$ that computes values of a “hidden” function $J_{C}\colon\{0,1\}^{m}\to G$ and the problem of learning $C$.
Problem 8.10.
Investigate computational complexity of learning $C$ from a hidden function $J_{C}$ for different (classes of) finite/infinite groups.
8.5. Spherical equations over infinite groups: decidability
For what (infinite) groups $G$ is the Diophantine problem for spherical equations decidable, but the problem for constrained spherical equations is not? Obviously any instance of ${\mathbf{CISE}}$ with a finite search space (e.g. when $|G|<\infty$ or when each variable can attain finitely many values) can be solved by enumerating all possible solutions. If $|G|=\infty$ and $Z_{j}$ are allowed to be infinite, then ${\mathbf{CISE}}$ can become undecidable. This question depends on how we allow to constrain variables, e.g.
  • sets $Z_{j}$ can be finitely generated subgroups of $G$,
  • or (more generally) sets $Z_{j}$ can be defined by rational subsets of $G$.
Problem 8.11.
For what (infinite) groups $G$ the Diophantine problem for spherical equations is decidable, but spherical equations with constraints of the form $z_{j}\in Z_{j}$, where $Z_{j}$ is a finitely generated subgroup of $G$ is not decidable.
Obvious candidates are groups containing a subgroup $H$ with undecidable membership problem, such as partially commutative groups. We believe that for these groups the Diophantine problem for spherical equations belongs to ${\mathbf{NP}}$, but the problem for the constrained conjugacy equations
$\left\{\begin{array}[]{l}z^{-1}cz=c\\ z\in H,\end{array}\right.$
is not decidable. We also formulate a similar problem for rationally constrained spherical equations.
Problem 8.12.
For what (infinite) groups $G$ the Diophantine problem for spherical equations is decidable, but spherical equations with constraints of the form $z_{j}\in Z_{j}$, where $Z_{j}$ is a rational subset of $G$, is not decidable.
Another class of groups with undecidable membership problem is the braid groups. For the braid groups it is not known how to approach even unconstrained spherical equations.
Problem 8.13.
Is the Diophantine problem for spherical (quadratic) equations over braid groups decidable?
8.6. Spherical equations over infinite groups: universality
Recall that a set of functions $H=\{h\colon D\to R\}$ is universal if for any distinct $x,y\in D$
$\Pr[h(x)=h(y)]\leq\tfrac{1}{|R|},$
where the probability is taken over a uniformly chosen $h\in H$. This condition can be relaxed by allowing $\Pr[h(x)=h(y)]$ to be $O(\tfrac{1}{|R|})$. Notice that in this paper for $R$ we have used certain classes of finite groups only (namely, $\{{\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}\}$ and $\{S_{n}\}$). Is it possible to construct a family of efficient universal functions using a single infinite group?
Problem 8.14.
Design a family of universal $0/1$-spherical functions or $J_{C}$ functions with the ranges $R$ being a part of the same infinite group.
8.7. Spherical functions: public-key encryption
Problem 8.15.
Is it possible to design a public-key encryption scheme which security is based on computational hardness of solving constrained spherical equations (or another related group-theoretic problem)?

### v2

@@
 Abstract.
-In this paper we analyze computational properties of the Diophantine problem (and its search variant) for spherical equations $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ (and its variants) over the class of finite metabelian groups $G_{p,n}={\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$, where $n\in{\mathbb{N}}$ and $p$ is prime.
+In this paper we analyze computational properties of the Diophantine problem (and its search variant) for spherical equations $\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=1$ (and their variants) over the class of finite metabelian groups $G_{p,n}={\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$, where $n\in{\mathbb{N}}$ and $p$ is prime.
 We prove that the problem of finding solutions for certain constrained spherical equations is computationally hard on average (assuming that some lattice approximation problem is hard in the worst case).
-Keywords. Spherical equations, finite groups, semidirect products, metabelian groups, average case complexity, hash function family, group-based cryptography.
+Key words and phrases:
+Spherical equations, finite groups, semidirect products, metabelian groups, average case complexity, hash function family, group-based cryptography.
 2020 Mathematics Subject Classification. 20F16, 20F10, 68W30.
@@
 For a group $G$, an equation over $G$ with variables in $Z$ is an equality of the form $W=1$, where $W\in F\ast G$.
-If $W=z_{i_{1}}g_{1}\cdots z_{i_{k}}g_{k}$, with $z_{i_{j}}\in Z$ and $g_{j}\in G$, then we refer to $\{z_{i_{1}},\ldots,z_{i_{k}}\}$ as the set of variables and to $\{g_{1},\ldots,g_{k}\}$ as the set of constants (or coefficients) of $W$.
+If $W=z_{i_{1}}^{\varepsilon_{1}}g_{1}\cdots z_{i_{k}}^{\varepsilon_{k}}g_{k}$, with $z_{i_{j}}\in Z$, $\varepsilon_{j}=\pm 1$, and $g_{j}\in G$, then we refer to $\{z_{i_{1}},\ldots,z_{i_{k}}\}$ as the set of variables and to $\{g_{1},\ldots,g_{k}\}$ as the set of constants (or coefficients) of $W$.
 We occasionally write $W(z_{1},\ldots,z_{k})$ or $W(z_{1},\ldots,z_{k};g_{1},\ldots,g_{k})$ to indicate that the variables in $W$ are precisely $z_{1},\ldots,z_{k}$ and (in the latter case) the constants are precisely $g_{1},\ldots,g_{k}$.
@@
 In this paper we work with a set of the form $Z=Z_{1}\times\dots\times Z_{n}$, in which case every variable $z_{i}$ is constrained individually by $Z_{i}$.
-Constrained equations in groups have received very little attention; only certain equations in free groups have been studied.
-For instance, in [9] V.
-Diekert demonstrated decidability and ${\mathbf{PSPACE}}$-hardness of the following problem: given an arbitrary equation
-$w(x_{1},\dots,x_{n},a_{1},\dots,a_{m})=1$
-in a free group $F_{m}$ and a list $H_{1},\dots,H_{n}$ of regular subsets of $F_{m}$, decide whether there exists a solution of this equation satisfying $x_{i}\in H_{i}$.
-V.
-Durnev showed in [11] that for a free group $F=F(a,b)$ there exists an equation of type
-$w(x^{k},x_{1},\dots,x_{n})=[a,b]$
-for which there is no algorithm to recognize if, for a given $k$, it has a solution satisfying $x_{1},\dots,x_{n}\in[F_{2},F_{2}]$.
+Recently, constrained equations in groups have attracted significant attention, with most research focusing on free groups, hyperbolic groups, right-angled Artin groups, virtually abelian groups, and certain classes of finite groups.
+For a survey of recent results in this area see [7].
 1.3. Spherical equations
@@
 It is a well known consequence of the classification of compact surfaces that any quadratic equation over $G$ is equivalent, via an automorphism $\phi$, computable in time $O(|W|^{2})$, to an equation in exactly one of the following three standard forms (see [8, 16]):
-(1) $\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $m\geq 1,$
-(2) $\prod_{i=1}^{g}[x_{i},y_{i}]\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0,$
-(3) $\prod_{i=1}^{g}x_{i}^{2}\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0.$
+$\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $m\geq 1,$ (1)
+$\prod_{i=1}^{g}[x_{i},y_{i}]\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0,$ (2)
+$\prod_{i=1}^{g}x_{i}^{2}\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}$ $=1$ $g\geq 1,m\geq 0.$ (3)
 The number $g$ is the genus of the equation, and both $g$ and $m$ (the number of constants) are invariants.
@@
 It is easy to show that an equation of the form
-(4) $\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}=c$
+$\prod_{j=1}^{m}z_{j}^{-1}c_{j}z_{j}=c$ (4)
 is equivalent to a spherical equation. We say that an equation (4) is an inhomogeneous form of a spherical equation.
@@
   • Computing the multiplicative inverse of a unit modulo $p$ can be done in $O(\lceil\log_{2}(p)\rceil^{2})$ time using the extended Euclidean algorithm.
-2. Preliminaries: knapsack problems in groups
+2. Preliminaries: subset sum problem in groups
 Here we review several useful definitions of discrete optimization in groups.
@@
 The subset sum problem ${\mathbf{SSP}}(G,X)$: Given $g_{1},\ldots,g_{k},g\in G$ decide if
-(5) $g=g_{1}^{\varepsilon_{1}}\cdots g_{k}^{\varepsilon_{k}}$
+$g=g_{1}^{\varepsilon_{1}}\cdots g_{k}^{\varepsilon_{k}}$ (5)
 for some $\varepsilon_{1},\ldots,\varepsilon_{k}\in\{0,1\}$.
-By [31, Proposition 2.5], computational properties of ${\mathbf{SSP}}$ do not depend on the choice of a finite generating set $X$ and, hence, the problem can be abbreviated as ${\mathbf{SSP}}(G)$.
+By [30, Proposition 2.5], computational properties of ${\mathbf{SSP}}$ do not depend on the choice of a finite generating set $X$ and, hence, the problem can be abbreviated as ${\mathbf{SSP}}(G)$.
 The same paper, [30], provides a variety of examples of groups with ${\mathbf{NP}}$-complete (or polynomial time) subset sum problems.
 Consider the infinitely generated group ${\mathbb{Z}}_{3}^{\omega}$ whose elements can be formally viewed as functions $f\colon{\mathbb{N}}\to{\mathbb{Z}}_{3}$ with finite support.
-For algorithmic purposes, we assume that elements of ${\mathbb{Z}}_{3}^{\omega}$ are encoded by finite ternary strings (as in [35, Section 4]).
-[36, Proposition 2.1] proves that ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{\omega})$ is ${\mathbf{NP}}$-complete, which can be reformulated as follows.
-Proposition 2.1 ([36, Proposition 2.1]).
+For algorithmic purposes, we assume that elements of ${\mathbb{Z}}_{3}^{\omega}$ are encoded by finite ternary strings (as in [34, Section 4]).
+[35, Proposition 2.1] proves that ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{\omega})$ is ${\mathbf{NP}}$-complete, which can be reformulated as follows.
+Proposition 1 ([35, Proposition 2.1]).
 For every $m\geq 3$, ${\mathbf{SSP}}$ is ${\mathbf{NP}}$-complete for the class of finite groups $\{{\mathbb{Z}}_{m}^{n}\}_{n=1}^{\infty}$.
@@
 Short integer solution (${\mathbf{SIS}}$) problem. For a given matrix $A\in{\mathbb{Z}}_{p}^{n\times m}$, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ satisfying
-(7) $\left\{\begin{array}[]{l}A{\overline{x}}\equiv_{p}{\overline{0}},\\ {\overline{x}}\mbox{ is short and nontrivial},\\ \end{array}\right.$
+$\left\{\begin{array}[]{l}A{\overline{x}}\equiv_{p}{\overline{0}},\\ {\overline{x}}\mbox{ is short and nontrivial},\\ \end{array}\right.$ (7)
 assuming that a solution exists.
@@
 Randomized ${\mathbf{SIS}}$ problem.
-For a matrix $A\leftarrow{\mathbb{Z}}_{p}^{n\times m}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{0}}$, and ${\overline{x}}$ is short and nontrivial.
-Theorem 3.1 (Goldreich, Goldwasser, Halevi, [16, Theorem 1], cf. Ajtai, [3, Theorem 1]).
+For a matrix $A\in{\mathbb{Z}}_{p}^{n\times m}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{0}}$, and ${\overline{x}}$ is short and nontrivial.
+Theorem 2 (Goldreich, Goldwasser, Halevi, [15, Theorem 1], cf. Ajtai, [3, Theorem 1]).
 Suppose that a PPT algorithm ${\mathcal{A}}$ solves the randomized ${\mathbf{SIS}}_{\{0,1\}}$ problem (or ${\mathbf{SIS}}_{\{-1,0,1\}}$ problem) with parameters $n,m,p$ satisfying
-(8) $n\log(p)<m<\tfrac{p}{2n^{4}}\ \mbox{ and }\ p=O(n^{c})\mbox{ for some }c>0,$
+$n\log(p)<m<\tfrac{p}{2n^{4}}\ \mbox{ and }\ p=O(n^{c})\mbox{ for some }c>0,$ (8)
 with probability at least $n^{-c_{0}}$ for some fixed constant $c_{0}>0$, where the probability is taken over the choice of the instance as well as the coin-tosses of ${\mathcal{A}}$.
@@
 Randomized ${\mathbf{ISIS}}$ problem.
-For a matrix $A\leftarrow{\mathbb{Z}}_{p}^{n\times m}$ and a vector ${\overline{y}}\in{\mathbb{Z}}_{p}^{n}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{y}}$, and ${\overline{x}}$ is short.
-Corollary 3.2.
-The statement of Theorem 3.1 holds for the randomized ${\mathbf{ISIS}}_{\{0,1\}}$ problem.
-Proof.
+For a matrix $A\in{\mathbb{Z}}_{p}^{n\times m}$ and a vector ${\overline{y}}\in{\mathbb{Z}}_{p}^{n}$ sampled uniformly randomly, find ${\overline{x}}\in{\mathbb{Z}}_{p}^{m}$ such that $A{\overline{x}}={\overline{y}}$, and ${\overline{x}}$ is short.
+Corollary 3.
+The statement of Theorem 2 holds for the randomized ${\mathbf{ISIS}}_{\{0,1\}}$ problem.
+Proof 3.1.
 Indeed, suppose that there is a PPT algorithm ${\mathcal{A}}$ that solves ${\mathbf{ISIS}}_{\{0,1\}}$ problem with probability at least $n^{-c_{0}}$.
@@
     solve $(A^{\prime},-{\overline{v}}_{i})$ using ${\mathcal{A}}$.
-This gives a PPT algorithm that solves ${\mathbf{SIS}}_{\{0,1\}}$ with probability at least $n^{-c_{0}}$. ∎
-4. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
+This gives a PPT algorithm that solves ${\mathbf{SIS}}_{\{0,1\}}$ with probability at least $n^{-c_{0}}$.
+4. Spherical equations over $G_{p,n}$
 In this section we analyse complexity of solving a spherical equation (1) over $G_{p,n}$, with constants $c_{i}=({\overline{c}}_{i},\beta_{i})\in G_{p,n}$ and unknowns $z_{1},\dots,z_{m}$.
-Lemma 4.1.
+Lemma 4.
 […]
@@
     $\sum_{i=1}^{m}B_{i}\alpha_{i}^{-1}((\beta_{i}-1){\overline{z}}_{i}+{\overline{c}}_{i})\equiv_{p}{\overline{0}}$, where $B_{i}=\beta_{1}\cdots\beta_{i-1}$.
-Proof.
-Straightforward verification. ∎
-4.1. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$: generic-case
-Here, we investigate generic-case hardness for spherical equations, i.e., hardness of a “typical” equation, see [32, Chapter 10] for basic definitions of generic-case complexity.
-Proposition 4.2.
+Proof 4.1.
+Straightforward verification.
+4.1. Spherical equations over $G_{p,n}$: generic-case
+Here, we investigate generic-case hardness for spherical equations, i.e., hardness of a “typical” equation, see [31, Chapter 10] for basic definitions of generic-case complexity.
+Proposition 5.
 If $\beta_{i}\neq 1$ for some $i$, then (1) has a solution and a solution can be found in polynomial time.
-Proof.
+Proof 4.2.
 It takes nearly linear time to compute $B_{1},\dots,B_{m}$ and check the condition (S1). If $\beta_{i}\neq 1$, then the following assignment:
@@
   • ${\overline{z}}_{i}\ \ \equiv_{p}\ \ \tfrac{1}{B_{i}(\beta_{i}-1)}\sum_{j=1}^{m}B_{j}{\overline{c}}_{j}$
-is a solution that can be computed in polynomial time. ∎
+is a solution that can be computed in polynomial time.
 Next, we claim that the property $\exists i\ \beta_{i}\neq 1$ is strongly generic, i.e., a typical equation satisfies this property.
@@
 $|{\mathcal{I}}_{s}|$ $=\sum_{p,n,m\leq s}(p^{n}\cdot(p-1))^{m}.$
-Lemma 4.3.
+Lemma 6.
 $\Pr\left\{\,(E_{m},G_{p,n})\in{\mathcal{I}}_{s}\;\middle|\;m\geq s/2\,\right\}\to 1$ exponentially fast as $s\to\infty$.
-Proof.
+Proof 4.3.
 For any fixed $p\geq 2,n\geq 1$ we have
@@
 $\frac{V_{s/2}}{V_{s}}=\frac{(p^{n}\cdot(p-1))^{s/2+1}-1}{(p^{n}\cdot(p-1))^{s+1}-1}\leq\frac{1}{(p^{n}\cdot(p-1))^{s/2}}\leq\frac{1}{2^{s/2}}$
-that converges to 00 exponentially fast. The obtained bound $\frac{1}{2^{s/2}}$ does not depend on $p$ or $n$. Therefore, the bound works on the whole ${\mathcal{I}}_{s}$. ∎
-Lemma 4.4.
+that converges to 00 exponentially fast. The obtained bound $\frac{1}{2^{s/2}}$ does not depend on $p$ or $n$. Therefore, the bound works on the whole ${\mathcal{I}}_{s}$.
+Lemma 7.
 $\Pr\left\{\,(E_{m},G_{p,n})\in{\mathcal{I}}_{s}\;\middle|\;\exists i\ \beta_{i}\neq 1\,\right\}\to 1$ exponentially fast as $s\to\infty$.
-Proof.
-By Lemma 4.3, we may assume that $m\geq s/2$, in which case
+Proof 4.4.
+By Lemma 6, we may assume that $m\geq s/2$, in which case
 $\Pr\left\{\,(E_{m},G_{p,n})\in{\mathcal{I}}_{s}\;\middle|\;\exists i\ \beta_{i}\neq 1\,\right\}\geq 1-\frac{1}{(p-1)^{s/2}},$
-which converges to $1$ exponentially fast as $s\to\infty$. ∎
-Corollary 4.5.
+which converges to $1$ exponentially fast as $s\to\infty$.
+Corollary 8.
 The Diophantine (decision) problem for spherical equations over $\{G_{p,n}\}$ is decidable in strongly generically linear time.
 The corresponding search problem can be solved in strongly generically polynomial time.
-Proof.
-Follows from Proposition 4.2 and Lemma 4.4
+Proof 4.5.
+Follows from Proposition 5 and Lemma 7
 Since spherical equations in which at least one coefficient $c_{i}=({\overline{c}}_{i},\beta_{i})$ satisfies $\beta_{i}\neq 1$ are computationally easy, we put a restriction on the coefficients that we use in spherical equations.
 Define the set
-(9) $C_{p,n}=\left\{\,({\overline{c}},1)\;\middle|\;{\overline{c}}\in{\mathbb{Z}}_{p}^{n}\,\right\}.$
-From now on we assume that all $c_{i}\in C_{p,n}$. In that case condition (S1) of Lemma 4.1 is trivially satisfied and condition (S2) translates into the following:
-(10) $\exists\alpha_{i}\in{\mathbb{Z}}_{p}^{\ast}\ \ \mbox{ s.t. }\ \sum_{i=1}^{m}\alpha_{i}^{-1}{\overline{y}}_{i}\equiv_{p}{\overline{0}}.$
+$C_{p,n}=\left\{\,({\overline{c}},1)\;\middle|\;{\overline{c}}\in{\mathbb{Z}}_{p}^{n}\,\right\}.$ (9)
+From now on we assume that all $c_{i}\in C_{p,n}$. In that case condition (S1) of Lemma 4 is trivially satisfied and condition (S2) translates into the following:
+$\exists\alpha_{i}\in{\mathbb{Z}}_{p}^{\ast}\ \ \mbox{ s.t. }\ \sum_{i=1}^{m}\alpha_{i}^{-1}{\overline{y}}_{i}\equiv_{p}{\overline{0}}.$ (10)
 The obtained condition defines a homogeneous system of linear congruences that does not allow zero values for unknowns $\alpha_{i}$, which makes the problem nontrivial (we conjecture it is ${\mathbf{NP}}$-hard).
-4.2. Spherical equations over ${\mathbb{Z}}_{p,n}\rtimes{\mathbb{Z}}_{p}^{\ast}$: worst case
+4.2. Spherical equations over $G_{p,n}$: worst case
 Here we investigate the worst case complexity for spherical equations. For ${\overline{w}}\in{\mathbb{Z}}_{3}^{n}$ define
@@
 For $I$ define ${\overline{v}}^{\prime}={\overline{v}}+\sum{\overline{v}}_{i}$ and the spherical equation $E_{I}$ as
-(11) $\prod_{i}^{m}z_{i}^{-1}c_{{\overline{v}}_{i}}z_{i}=c_{{\overline{v}}^{\prime}}.$
-Proposition 4.6.
+$\prod_{i}^{m}z_{i}^{-1}c_{{\overline{v}}_{i}}z_{i}=c_{{\overline{v}}^{\prime}}.$ (11)
+Proposition 9.
 $I$ is a positive instance of ${\mathbf{SSP}}$Leftrightarrow\ \ $E_{I}$ has a solution.
-Proof.
+Proof 4.6.
 By design we have
@@
 […]
-Corollary 4.7.
+Corollary 10.
 The Diophantine problem for spherical equations over ${\mathbb{Z}}_{3,n}\rtimes{\mathbb{Z}}_{3}^{\ast}$ is ${\mathbf{NP}}$-complete.
-Proof.
-By Proposition 4.6, $I\to E_{I}$ is a many-one polynomial-time reduction of ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{n})$ to spherical equations over ${\mathbf{SSP}}({\mathbb{Z}}_{3,n}\rtimes{\mathbb{Z}}_{3}^{\ast})$.
-By Proposition 2.1, ${\mathbf{SSP}}$ is ${\mathbf{NP}}$-complete for groups $\{{\mathbb{Z}}_{3}^{n}\}_{n=1}^{\infty}$.
+Proof 4.7.
+By Proposition 9, $I\to E_{I}$ is a many-one polynomial-time reduction of ${\mathbf{SSP}}({\mathbb{Z}}_{3}^{n})$ to spherical equations over ${\mathbf{SSP}}({\mathbb{Z}}_{3,n}\rtimes{\mathbb{Z}}_{3}^{\ast})$.
+By Proposition 1, ${\mathbf{SSP}}$ is ${\mathbf{NP}}$-complete for groups $\{{\mathbb{Z}}_{3}^{n}\}_{n=1}^{\infty}$.
 Hence the result.
-∎
 5. Constrained spherical equations: average-case hardness
-In this section we discuss spherical equations over the groups $G_{p,n}$ with constraints on values of $z_{i}$’s.
+In this section we discuss spherical equations over the class of groups ${\mathcal{G}}=\{G_{p,n}\}$ with constraints on values of $z_{i}$’s.
 In full generality the problem can be formulated as follows.
@@
 It is not hard to prove that ${\mathbf{CISE}}$ for groups $G_{p,n}$ is hard in the worst case.
-In fact, for $p=3$, the proof of Proposition 4.6 works as is, with constraints $Z_{i}=\{({\overline{0}},1),({\overline{0}},2)\}$ (so chosen sets $Z_{i}$ effectively define no constraints).
+In fact, for $p=3$, the proof of Proposition 9 works as is, with constraints $Z_{i}=\{({\overline{0}},1),({\overline{0}},2)\}$ (so chosen sets $Z_{i}$ effectively define no constraints).
 Using these constraints we can easily achieve the same result for any odd value of $p$.
-Theorem 5.1.
+Theorem 11.
 ${\mathbf{CISE}}$ is ${\mathbf{NP}}$-hard for groups $\{G_{p,n}\}_{n}$ for any fixed odd prime $p$.
-Proof.
-For a given instance ${\overline{v}}_{1},\dots,{\overline{v}}_{m},{\overline{v}}\in{\mathbb{Z}}_{p}^{n}$, call it $I$, of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$, as in the proof of Proposition 4.6, construct a constrained inhomogeneous spherical equation $E_{I}$ over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
+Proof 5.1.
+For a given instance ${\overline{v}}_{1},\dots,{\overline{v}}_{m},{\overline{v}}\in{\mathbb{Z}}_{p}^{n}$, call it $I$, of ${\mathbf{SSP}}({\mathbb{Z}}_{p}^{n})$, as in the proof of Proposition 9, construct a constrained inhomogeneous spherical equation $E_{I}$ over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$
 $\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{{\overline{v}}_{i}}z_{i}=c_{{\overline{v}}^{\prime}}\\ z_{i}=({\overline{0}},1)\mbox{ or }({\overline{0}},2^{-1})\end{array}\right.$
@@
 This gives a polynomial time reduction from ${\mathbf{SSP}}$ over groups $\{{\mathbb{Z}}_{p}^{n}\}_{n=1}^{\infty}$ to decidability of constrained inhomogeneous spherical equations over ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$.
-∎
 Our next goal is to demonstrate the average-case hardness of constrained spherical equations. To achieve this, we randomize equations as follows. First, we stratify parameters:
@@
 Then, for an arbitrary but fixed $n\in{\mathbb{N}}$, consider the set of coefficients $C_{p,n}$ defined in (9) and a (finite) set of constrained inhomogeneous spherical equations
-(12) $\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{n}\\ z_{i}=({\overline{0}},1)\mbox{ or }({\overline{0}},2^{-1}).\end{array}\right.$
+$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{p,n}\\ z_{i}=({\overline{0}},1)\mbox{ or }({\overline{0}},2^{-1}).\end{array}\right.$ (12)
 The problem to find a solution of (12) is called ${\mathbf{CISE}}_{\{1,2\}}$. Also, we want to study slightly relaxed equations
-(13) $\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{n}\\ z_{i}=({\overline{0}},1),({\overline{0}},2^{-1}),\mbox{ or }({\overline{0}},3^{-1}).\end{array}\right.$
+$\left\{\begin{array}[]{l}\prod_{i=1}^{m}z_{i}^{-1}c_{i}z_{i}=c\ \ \mbox{ with }c_{i},c\in C_{p,n}\\ z_{i}=({\overline{0}},1),({\overline{0}},2^{-1}),\mbox{ or }({\overline{0}},3^{-1}).\end{array}\right.$ (13)
 The problem to find a solution of (13) is called ${\mathbf{CISE}}_{\{1,2,3\}}$.
 Randomized constrained inhomogeneous spherical equation problem. Find a solution for a uniformly distributed system (12) (or (13)).
-Theorem 5.2.
+Theorem 12.
 Suppose that a PPT algorithm ${\mathcal{A}}$ solves the randomized ${\mathbf{CISE}}_{\{1,2\}}$ problem (or ${\mathbf{CISE}}_{\{1,2,3\}}$ problem) with parameters $n,m,p$ satisfying (8) with probability at least $n^{-c_{0}}$ for some fixed constant $c_{0}>0$, where the probability is taken over the choice of the instance as well as the coin-tosses of ${\mathcal{A}}$.
@@
 $1-2^{-n}$).
-Proof.
+Proof 5.2.
 Consider an arbitrary instance of $(A,{\overline{y}})$ of ${\mathbf{ISIS}}_{\{0,1\}}$.
@@
 […]
-This establishes a one-to-one correspondence between ${\mathbf{ISIS}}_{\{0,1\}}$ and ${\mathbf{CISE}}_{\{1,2\}}$ that translates the uniform distribution for the instances of ${\mathbf{ISIS}}_{\{0,1\}}$ to the uniform distribution for the instances of ${\mathbf{CISE}}_{\{1,2\}}$ (the element ${\overline{v}}^{\prime}\in C_{n}$ is uniformly distributed because ${\overline{y}}$ uniformly distributed).
+This establishes a one-to-one correspondence between ${\mathbf{ISIS}}_{\{0,1\}}$ and ${\mathbf{CISE}}_{\{1,2\}}$ that translates the uniform distribution for the instances of ${\mathbf{ISIS}}_{\{0,1\}}$ to the uniform distribution for the instances of ${\mathbf{CISE}}_{\{1,2\}}$ (the element ${\overline{v}}^{\prime}\in C_{p,n}$ is uniformly distributed because ${\overline{y}}$ uniformly distributed).
 A similar reduction can be used to establish a one-to-one correspondence between ${\mathbf{SIS}}_{\{-1,0,1\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$.
-∎
-Remark 1.
-The one-to-one correspondence established in the proof of Theorem 5.2 also implies that uniformly chosen instances of ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$ have solutions with probability approaching $1$ as $n\to\infty$.
-Remark 2.
+Remark 5.3.
+The one-to-one correspondence established in the proof of Theorem 12 also implies that uniformly chosen instances of ${\mathbf{CISE}}_{\{1,2\}}$ and ${\mathbf{CISE}}_{\{1,2,3\}}$ have solutions with probability approaching $1$ as $n\to\infty$.
+Remark 5.4.
 Similar results can be proven for the class of symmetric groups $\{S_{n}\}_{n=1}^{\infty}$, or any other class of groups that contain the groups $G_{p,n}$ in a “compact” way.
@@
 ${\mathcal{H}}_{n}=\left\{\,H_{\overline{c}}\;\middle|\;c_{1},\dots,c_{m}\in C_{p,n}\,\right\}\mbox{ and }{\mathcal{H}}=\bigcup_{n=1}^{\infty}{\mathcal{H}}_{n}.$
-Proposition 6.1.
+Proposition 13.
 If ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ is hard in the worst case, then ${\mathcal{H}}$ is a one-way function family.
-Proof.
+Proof 6.1.
 To prove that ${\mathcal{H}}$ is a one-way function family it is sufficient to show that for every PPT algorithm ${\mathcal{A}}$ the sequence of values
@@
 Therefore, if it is not true that $P_{\mathcal{A}}(n)$ is $o(n^{-d})$ for some PPT algorithm ${\mathcal{A}}$ and some $d>0$, then $P_{\mathcal{A}}(n)\geq c\cdot n^{-d}$ (for some $c>0$) satisfied for infinitely many indices $n$.
-By Theorem 5.2 that means that there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6).
-∎
-Proposition 6.2.
+By Theorem 12 that means that there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6).
+Proposition 14.
 If ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ is hard in the worst case, then ${\mathcal{H}}$ is a collision-free function family.
-Proof.
+Proof 6.2.
 To prove that ${\mathcal{H}}$ is a collision-free hash function family it is sufficient to show that for every PPT algorithm ${\mathcal{A}}$ the sequence of values
@@
 Thus, if we can efficiently find collisions for uniformly distributed $0/1$-spherical equations $H_{\overline{c}}\in{\mathcal{H}}_{n}$ for infinitely many indices $n$, then we can efficiently find solutions for uniformly distributed ${\mathbf{CISE}}_{\{1,2,3\}}$.
-Then, by Theorem 5.2, there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6).
+Then, by Theorem 12, there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6).
 Contradiction.
-∎
 7. The acyclic graph word problem: average-case hardness
@@
 Figure 1. An instance of ${\mathbf{AGWP}}$ corresponding to an instance of ${\mathbf{CISE}}$ (assuming that $Z_{j}=\{g_{1},g_{2}\}$).
-Proposition 7.1.
+Proposition 15.
 The instance of ${\mathbf{CISE}}$ is positive if and only if the instance of ${\mathbf{AGWP}}$ is positive.
 The reduction described above induces a system of probability measures $\{\mu_{n}\}_{n\in{\mathbb{N}}}$ on instances of ${\mathbf{AGWP}}$ over the class of groups ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$.
-In fact, each $\mu_{n}$ gives uniform distribution on a subset of instances of ${\mathbf{AGWP}}$ of type shown in Figure 7.1.
+In fact, each $\mu_{n}$ gives uniform distribution on a subset of instances of ${\mathbf{AGWP}}$ of type shown in Figure 15.
 That establishes a one-to one correspondence between uniformly randomized ${\mathbf{CISE}}_{\{1,2\}}$ described in Section 5 and ${\mathbf{AGWP}}$ endowed with $\{\mu_{n}\}_{n\in{\mathbb{N}}}$, which implies average-case hardness of ${\mathbf{AGWP}}$ (again, assuming (6)).
@@
 8.1. Spherical equations over finite groups: constrained versus unconstrained
-Problem 8.1.
+Problem 16.
 Does there exist a class of finite groups ${\mathcal{G}}$ satisfying the following conditions:
@@
 It was proved in [29] that the Diophantine problem for spherical equations over ${\mathcal{G}}=\{\operatorname{{GL}}(2,p)\}_{p\in\mathbb{P}}$ can be solved efficiently.
-Problem 8.2.
-Does $\{\operatorname{{GL}}(2,p)\}_{p\in\mathbb{P}}$ satisfy the second condition of Problem 8.1?
-Problem 8.2 can be generalized as follows.
-Problem 8.3.
-Does any family of classical finite groups (such as $\operatorname{{SL}}(n,p)$, $\operatorname{{PSL}}(n,p)$, etc., see [24]) satisfy conditions of Problem 8.1?
+Problem 17.
+Does $\{\operatorname{{GL}}(2,p)\}_{p\in\mathbb{P}}$ satisfy the second condition of Problem 16?
+Problem 17 can be generalized as follows.
+Problem 18.
+Does any family of classical finite groups (such as $\operatorname{{SL}}(n,p)$, $\operatorname{{PSL}}(n,p)$, etc., see [23]) satisfy conditions of Problem 16?
 8.2. Constrained spherical equations: foundation of average-case hardness
@@
 Notice that existence of uniform self-reductions for ${\mathbf{ISIS}}$ (like for the discrete logarithm problem) appears to be implausible.
-Problem 8.4.
+Problem 19.
 Design a class of randomized self-reductions $\Phi$ for ${\mathbf{ISIS}}_{\{0,1\}}$ and use them to prove the following.
@@
 Similarly, it would be interesting to investigate self-reductions for ${\mathbf{CISE}}$ over some groups ${\mathcal{G}}$ and use them to demonstrate that the worst-case hardness for ${\mathbf{CISE}}$ implies its average-case hardness.
-Problem 8.5.
+Problem 20.
 For a class of spherical equations over a class of groups ${\mathcal{G}}$, design a class of randomized self-reductions for ${\mathbf{CISE}}$ and use them to prove the following.
@@
 8.3. Constrained spherical equations: generalization
-As we have shown in Section 7, instances of ${\mathbf{CISE}}$ can be transformed into instances of ${\mathbf{AGWP}}$ of certain type (shown in Figure 7.1).
-Let us generalize the digraph shown in Figure 7.1 and modify it into a finite-state transducer.
+As we have shown in Section 7, instances of ${\mathbf{CISE}}$ can be transformed into instances of ${\mathbf{AGWP}}$ of certain type (shown in Figure 15).
+Let us generalize the digraph shown in Figure 15 and modify it into a finite-state transducer.
 For a table of elements $C=\{c_{ij}\}$ from $G$ (where $i=1,\dots,m$ and $j=0,1$) define a transducer $\Gamma_{C}$ shown in Figure 2.
@@
 Also, notice that the functions $J_{C}$ generalize hash function construction introduced in [39] and its numerous variations (e.g., [38, 6]), including those utilizing monoids.
-Problem 8.6.
-Investigate average-case hardness of computing $J_{C}^{-1}$.
-The most promising approach to Problem 8.6 appears to be via self-reductions similar to those discussed in Section 8.2.
+Problem 21.
+Investigate average-case hardness of computing $J_{C}^{-1}$ for $C$ sampled uniformly randomly.
+The most promising approach to Problem 21 appears to be via self-reductions similar to those discussed in Section 8.2.
 Furthermore, it might be easier to design self-reductions for functions $J_{C}$ because $J_{C}$ does not require pairs $c_{i0}$ and $c_{i1}$ to be conjugate.
-Problem 8.7.
+Problem 22.
 For a class of groups ${\mathcal{G}}=\{G_{n}\}$ design a randomized self-reduction for the problem of computing $J_{C}^{-1}$ satisfying the following: if $J_{C}^{-1}$ can be computed by a PPT algorithm on a non-negligible set of instances, then it can be computed by a PPT algorithm on every instance with overwhelming probability.
-Problem 8.8.
+Problem 23.
 For a class of groups ${\mathcal{G}}$ investigate cryptographic properties for the family of functions $\{J_{C}\}_{C\in G^{2\times m}}$.
@@
 To learn the hidden function means to find the secret element ${\overline{c}}$ using a number of sampled pairs $({\overline{b}},g)$.
-Problem 8.9.
+Problem 24.
 Investigate computational complexity of learning ${\overline{c}}$ from a hidden function $H_{\overline{c}}$ for different (classes of) finite/infinite groups.
 In a similar way, we can consider a device $D$ that computes values of a “hidden” function $J_{C}\colon\{0,1\}^{m}\to G$ and the problem of learning $C$.
-Problem 8.10.
+Problem 25.
 Investigate computational complexity of learning $C$ from a hidden function $J_{C}$ for different (classes of) finite/infinite groups.
@@
   • or (more generally) sets $Z_{j}$ can be defined by rational subsets of $G$.
-Problem 8.11.
+Problem 26.
 For what (infinite) groups $G$ the Diophantine problem for spherical equations is decidable, but spherical equations with constraints of the form $z_{j}\in Z_{j}$, where $Z_{j}$ is a finitely generated subgroup of $G$ is not decidable.
@@
 is not decidable. We also formulate a similar problem for rationally constrained spherical equations.
-Problem 8.12.
+Problem 27.
 For what (infinite) groups $G$ the Diophantine problem for spherical equations is decidable, but spherical equations with constraints of the form $z_{j}\in Z_{j}$, where $Z_{j}$ is a rational subset of $G$, is not decidable.
 Another class of groups with undecidable membership problem is the braid groups. For the braid groups it is not known how to approach even unconstrained spherical equations.
-Problem 8.13.
+Problem 28.
 Is the Diophantine problem for spherical (quadratic) equations over braid groups decidable?
@@
 Is it possible to construct a family of efficient universal functions using a single infinite group?
-Problem 8.14.
+Problem 29.
 Design a family of universal $0/1$-spherical functions or $J_{C}$ functions with the ranges $R$ being a part of the same infinite group.
 8.7. Spherical functions: public-key encryption
-Problem 8.15.
+Problem 30.
 Is it possible to design a public-key encryption scheme which security is based on computational hardness of solving constrained spherical equations (or another related group-theoretic problem)?